OpenStudy (anonymous):

Consider the sequence 2, 6, 18, 54… Let n = represent the term number in the sequence. Let A(n) = the value of the nth term of the sequence. 4. What is the common ratio of the sequence?

OpenStudy (anonymous):

sorry Im not good with patterns at all

OpenStudy (anonymous):

A(n) = a.r^n= 2.3^n ratio r is= 6/2=18/6 = 3

OpenStudy (anonymous):

Cool can you help me with 2 more?

OpenStudy (anonymous):

sure..say d qstn

OpenStudy (anonymous):

5. Complete each statement a. A(1) = 2 = 2 x 3^? b. A(2) = 6 = 2 x 3 = 2 x 3^? c. A(3) = 18 = 2 x 3 x 3 = 2 x 3^? d. A(4) = 54 = 2 x 3 x 3 x 3 = 2 x 3^? Here it is @ritez

OpenStudy (anonymous):

a. 0 b. 1 c. 2 d. 3

OpenStudy (anonymous):

6. What is the relationship between the exponent and the base 3 and the value of n?

OpenStudy (anonymous):

is it the next question?

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

Do you know it?

OpenStudy (anonymous):

i dont understand d qstn